OpenStudy (anonymous):

The graph of which quadratic equation is shown below? a u-shaped graph on a coordinate plane that is opening up and has a vertex of (1, −9) y = x2 − 2x − 8 y = x2 + 2x − 8 y = −x2 + 2x − 8 y = −x2 − 2x − 8

OpenStudy (debbieg):

When you look at the equation of a quadratic function, do you know how to determine if the parabola opens up, or opens down? That lets you eliminate two of these answers.

OpenStudy (debbieg):

Then, you COULD just check the point in the other two equations. Although, the more rigorous method would be to use the vertex for for a quadratic: y=a(x-h)^2-k plug in your given h and k, and multiply it out - see which equation you get.

OpenStudy (anonymous):

yes i know when a is positive the porabola is opened up but if it is negative it goes down and i also knowthat is the it's opened up it mans it's minimum and down is maximum that is all i know

OpenStudy (anonymous):

so that eliminates the two first answerss? @DebbieG

OpenStudy (anonymous):

@DebbieG sir are you there?

OpenStudy (debbieg):

Sorry, I stepped away from the computer. and I'm not a sir... lol. :) You are correct that a>0 means that the parabola opens up, and a<0 means that it opens down. but why do you think that the first 2 answers are eliminated? You said that your graph opens up... so a has to be....? positive, or negative?

OpenStudy (anonymous):

positive and sorry about calling you a sir the photo made me think it was you

OpenStudy (debbieg):

LOL I get that a lot. Those are my sons. :) OK, right - so a has to be positive. So which 2 answers does that rule out?

OpenStudy (anonymous):

the two last ones

OpenStudy (anonymous):

but in my graph the porabula looks like "u" not an "n" if you know what i'm trying to say

OpenStudy (debbieg):

Right - it opens upward. That means a>0. That means that your answer can't be C or D above, because they have a<0, right? So now you just need to decide if it is A or B.

OpenStudy (anonymous):

i'm sorry ma' am but the graph is going down so shouldn't we eliminate the top two (only asking questions to understand)

OpenStudy (anonymous):

oh nevermind sorry

OpenStudy (debbieg):

lol.. well, yes, that does change things. so it OPENS DOWN? Because above you said that it OPENED UP. OK, so are we back to it opens up??

OpenStudy (debbieg):

if it opens up, it's one of the first 2 equations. If it opens down, it's one of the last 2 equations.

OpenStudy (anonymous):

let me send you a picture ma' am

OpenStudy (debbieg):

ok

OpenStudy (anonymous):

this is what i have

OpenStudy (debbieg):

Yes, that opens up.

OpenStudy (debbieg):

so a>0

OpenStudy (anonymous):

ok

OpenStudy (debbieg):

So you have a vertex of (1, −9). Like I said above, there are a few different ways to proceed. If you knwo the "vertex form" of a parabola, you could use that and plug in the vertex: (1, −9) to the vertex form y=a(x-h)^2-k. You know that a=1, so you would just expand the (x-h)^2 and see which equation you get.

OpenStudy (anonymous):

so do you want me to find the vertex first?

OpenStudy (debbieg):

Or, you could "plug and chug" - trying the vertex in each of those 2 equations. See which graph it is on - if it's only one, then you know that's the one. I don't prefer that method though, because it is only BECAUSE this is a multiple choice problem that you can do it. it doesn't give you a rigorous understanding of the quadratic.

OpenStudy (debbieg):

Hmmm.... well, you HAVE the vertex. You could find the vertex of the equations given in the answer choices, that's another way to go about it.

OpenStudy (anonymous):

sorry lost connection but so you want me to plug in the 1 and -9?

OpenStudy (debbieg):

Well, I've given you several options. You know that it's one of the first two equations. there are a variety of ways to determine which one. Why don't you tell me what you think would work, and what your result is?

OpenStudy (anonymous):

I think we should use plug and chug since we do have the answer choises

OpenStudy (anonymous):

so should I plug and chug and tell you what I get? or do you want to do it together ma'am?

OpenStudy (anonymous):

i got b ma'am @DebbieG

OpenStudy (debbieg):

How did you get that?

OpenStudy (anonymous):

i asked my dad

OpenStudy (debbieg):

Does \(\large y = x^2 + 2x − 8\) have vertex (1, -9)? OK, no offense to your dad, but chances are good that you know more about quadratic equations than he does.

OpenStudy (debbieg):

Does \(\large y = x^2 + 2x − 8\) have vertex (1, -9)?

OpenStudy (debbieg):

Is (1, -9) a point ON THE GRAPH of \(\large y = x^2 + 2x − 8\)?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

i got 1 and -5

OpenStudy (debbieg):

You got (1, -5) as the vertex of \(\large y = x^2 + 2x − 8\)? Can you tell me how you got that?

OpenStudy (debbieg):

Or are you just saying that you tried x=1, and got that y=-5? THAT is correct. Which means that (1, -9) is NOT the vertex of this one, right?

OpenStudy (debbieg):

(1, -5) is not the vertex either; but it IS a point on the graph. However, (1, -9) is not even on the graph of this one.

OpenStudy (anonymous):

well i did x = -b/2a and then wrote x = -2/2(1) which is x = -2/2 = x= -1 and then did y = -1^2 + 2(-1) - 8 y = 1 - 2 - 8 y = 9

OpenStudy (anonymous):

oh i meant -1 and 9 not 1 and -5 oops

OpenStudy (anonymous):

so it is A

OpenStudy (anonymous):

because: x = -b/2a x = 2/2 x = 1 y = 1 - 2 - 8 y = 1-10 y = -9

OpenStudy (debbieg):

Ok, there is a big difference between (-1, 9) and (1, -5).... lol. But in any case, the equation you are looking for has a vertex of (1, -9). You are correct that the vertex of B has x=-1 (which is enough to rule out B as the answer). However, you made an error in finding the y coordinate of B: y = -1^2 + 2(-1) - 8 y = 1 - 2 - 8 <------ right here.... 1-2-8=-9 NOT =9 And YES you are correct - it is A. As your calculations show, THAT one has vertex (1, -9).

OpenStudy (anonymous):

ok thank you so much for your help without you i wouldn't have understood the problem or how to solve it and i feel like i learned a lot thank you so much for all your help and effort ma'am! :)